More Calculator Practice-Borrowing Capacity Use your calculator to solve the following problems. You can default to Excel if this is not possible. 1. You can afford to pay $6,000 a month for a 30-year mortgage. When interest rates are 6.125% how much can you borrow? Your answer: 2. If the loan represents 60% of the property's total value, how much is the property worth? Your answer: 3. How much equity do you require for the property? Your answer: 4. Suppose you want to use and interest only (IO) loan. Now how much could you borrow? Your answer: 5. How large of a property could you acquire now? Your answer :
More Calculator Practice-Medium Term Notes Use your calculator to solve the following problems. You can default to Excel if this is not possible. 1. How much would you pay per month to service (pay interest and principal) a 10-year, $3 million loan, fully amortizing loan? Your answer: 2. If the loan represents 55 % of the property's total value, how much the property worth? Your answer: 3. What would be your monthly payment on an IO loan? Your answer: 4. What would be your monthly payment if the loan had a 30-year amortization schedule? Click cell J14 for some help. The font is white. Either change it to another color or read the response in the formula bar. Your answer: 5. Assume you are able to pay the same monthly amount calculated in the first question above. How much could you borrow using a 30-year amortization schedule? Your answer What would be the most you could pay for a property using a 55% LTV (loan- -value, where the loan represents 55% of the total property value)? Your answer:
Answer #1

1. The money you can borrow is $987,474.64. Calculation given below:

Monthly interest rate=6.125%/12=0.510%

Total number of months you have to pay EMI=30*12=360 months

A PMT 6000 Number of months for loan tenure Monthly Interest Rate PV 360 0.510% ($987,474.64),

C PMT 6000 Number of months for loan tenure 360 Monthly Interest Rate PV 0.510% |=PV(B3,B2,B1,0) PV(rate, nper, pmt, [fv], [t

2. If the property value is X (Say) then

X*60%= 987,474.64 or X=$1645791.07

So, property value is $1,645,791.07

3. Equity needed to buy the property= (1,645,791.07- 987,474.64)=$6,58,316.4

4. Suppose, X is the amount I can borrow

So, X*0.510%=6000 or X=$1,176,471

So, I can borrow $1,176,471

5. Now the value of property I can buy= 1,176,471/0.6=$1960785